Question 5: (1 point) f(x) = + x - 18, find S (1) (a) 215 (b)-875 (c)-115 (d) 1/5 (e) -215
Answer #1

f(x) = 5x + x = x 3 + x fa) at x=1, Cf(!) x -1-8x28 +C = + (1) $ _ 118 (17 28 = + -108 = (- 8 6